a sodium vapor lamp emits photons with a wavelength of 5.89x10^-7m. What is the energy of these phothons
1
Expert's answer
2012-11-12T11:14:26-0500
E = h*c / λ;
E = 6,626 * 10^-34 J*s * 3*10^8 m/s/5.89x10^-7 m =3.37 *10^-19 J